# ![[Uncaptioned image]](x1.png)

### v1

@@ Abstract
Dimensionality Reduction (DR) techniques such as t-SNE and UMAP are popular for transforming complex datasets into simpler visual representations. However, while effective in uncovering general dataset patterns, these methods may introduce artifacts and suffer from interpretability issues. This paper presents DimVis, a visualization tool that employs supervised Explainable Boosting Machine (EBM) models (trained on user-selected data of interest) as an interpretation assistant for DR projections. Our tool facilitates high-dimensional data analysis by providing an interpretation of feature relevance in visual clusters through interactive exploration of UMAP projections. Specifically, DimVis uses a contrastive EBM model that is trained in real time to differentiate between the data inside and outside a cluster of interest. Taking advantage of the inherent explainable nature of the EBM, we then use this model to interpret the cluster itself via single and pairwise feature comparisons in a ranking based on the EBM model’s feature importance. The applicability and effectiveness of DimVis are demonstrated through two use cases involving real-world datasets, and we also discuss the limitations and potential directions for future research.
@@ 1 Introduction
Dimensionality Reduction (DR) is an essential component of data exploration, mainly transforming high-dimensional data into a more manageable, lower-dimensional form [VAT19]. This transformation aims to simplify complex datasets, improve computational efficiency, and retain essential information for analysis [SZS17]. Techniques such as t-SNE [MH08] and UMAP [MHM18] have gained popularity in DR for their ability to handle complex patterns and visualize high-dimensional data in distinct clusters. However, these nonlinear DR methods can make visualizations challenging to interpret due to suboptimal hyperparameters (e.g., [WVJ16]) and misleading distortions introduced in projections [HFA17, NA18].
To address these challenges, a plethora of visualization tools have been developed with the goal of inspecting DR projections interactively to assess the global and local quality of the 2D space [CMJK20, CMJ20, CKK24]. In addition to more complex visualization tools, recent works from Bibal et al. [BCDF21, BDF23] involved training interpretable supervised Machine Learning (ML) algorithms like decision trees to interpret DR projections contrastively. However, such methods suffer from scalability issues, with decision trees getting quickly deep and unparsable, if they overfit the data to achieve high fidelity. To overcome these limitations, Explainable Boosting Machine (EBM) [LCGH13], a tree-based, cyclic gradient boosting generalized additive model, is a powerful alternative that remains interpretable and scalable even with datasets with many features [CLG15].
Our proposed visualization tool, DimVis, employs an EBM model for unveiling the factors influencing cluster formation and data relationships in UMAP projections. It enables users to adjust UMAP hyperparameters and interactively inspect clusters to gain insights about global and local quality by retraining a supervised EBM model on user-selected data points and exploring single and pairs of features (ranked by EBM’s intrinsic feature importance). This comparative user-driven analysis of the selection against the remaining points is represented through intuitive and scalable visualizations such as bar charts, line plots, histograms, and heatmaps. The effectiveness of DimVis is demonstrated through a use case and a usage scenario with real-world healthcare datasets, showcasing the tool’s capability of analyzing high-dimensional datasets and enhancing the interpretability of the 2D space.
@@ 2 Related Work
Interpreting nonlinear DR projections is challenging due to the unknown high-dimensional feature positions that become obscure when transformed into a low-dimensional space. Many visualization tools focus on alleviating this problem (e.g., [PGO14, JPN15, FKM19]). Coimbra et al. [CMN16] introduced a method in which altering high-dimensional feature values generates 2D curved lines in projections to clarify feature behavior. Prolines [CD18] show the effect of individual feature changes in scatter plots by altering a single feature while maintaining others, simplifying high-dimensional data interpretation. Similarly, Turkay et al. [TFH11] and Yuan et al. [YRWG13] suggest a dual-view approach linking the original dimensions to 2D projections. This design enables users to observe how changes in one view impact the other, aiding in understanding the relationship between high- and low-dimensional features.
When interpreting DR projections through features, feature importance is vital [DB04]. Da Silva et al. [dSRM15] uses distance and variance metrics for high-dimensional feature ranking to detect key features defining clusters in the low-dimensional space. In contrast, DimVis uses the inherent EBM model’s feature importance.
More relevant to our work, t-viSNE [CMK20] supports the interactive assessment and interpretation of t-SNE projections. However, compared to DimVis, it is relatively complex with multiple coordinated views, and some of them are specifically designed for the t-SNE technique. DT-SNE [BDF23] is a simpler but static approach combining t-SNE with supervised decision trees for interpreting projections with features. Instead, IXVC [BCDF21] resolves this problem by facilitating users in interactively explaining projections with decision trees. However, decision trees, especially visualized as node-link diagrams, can become quickly uninterpretable, and the increase in the number of dataset features heavily affects their scalability. On the contrary, DimVis uses EBM – which is in most cases as powerful as ensemble learning models – and offers simple and scalable visual representations for exploring the most important features in user-picked data contrastively.
@@ 3 DimVis: System Overview and Use Case
DimVis is an open-source, web-based visualization tool [Dim23] developed in JavaScript and Python with libraries such as Plotly Dash [Plo13], Flask [Fla10], and InterpretML [NJKC19]. It employs a supervised EBM model to interpret visualizations generated with unsupervised DR techniques such as UMAP.
Tool overview and overall workflow. After selecting the dataset from the drop-down menu (shown in DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(a.1)), if the visualization result is not satisfactory, users can adjust two hyperparameters (located in DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(a.2)) until they find a visualization that meets their needs. Initially, only views indicated in (a.1), (a.2), and (b) are shown. To initiate the exploration and trigger the remaining views, a lasso selection of data points must be made in the UMAP view (b) containing a visual cluster perceived by the user as being of interest. Although the tool will work with any selection, regardless of whether it is a visual cluster or not, we assume that the selection includes a group of data points that encode an interesting visual feature of the dataset. As soon as a selection is made, an EBM model is trained in the background (more details below) and the remaining visualizations are presented. The model’s precision and recall are visible (in DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(a.3)), as well as single and pairwise feature importances extracted from the EBM model and visualized in a bar chart sorted in descending order from top to bottom (see DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(c)). On click, users can explore further a feature with a Score line plot (DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(d.1)) and a Density histogram (DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(d.2)), or two features with a Score heatmap.
@@
UMAP projection and hyperparameters. DimVis utilizes the UMAP algorithm [MHM18] that projects high-dimensional data into a 2D space, making it easier to visualize complex datasets. Users can interactively adjust UMAP’s hyperparameters – “Number of Neighbors” (Neighbors) and “Minimum Distance” (Min Dist) – to explore diverse projections (see Figure 1(a)). The projection in Figure 1(b) contains three distinct clusters (C1–C3), and no GT labels to simulate a typical unsupervised learning scenario.
@@
EBM’s performance metrics and feature importances. When a group of points is selected with the lasso functionality, the EBM gets retrained in a supervised manner where selected points are marked with class and the remaining points are marked with class . Making use of the inherent interpretability features of the EBM model, we rank the main features (single and pairs) that contribute to the separation of the cluster according to the trained model, using the Mean Absolute Score (Weighted). As shown in Figure 1(c), the most important single and pair of features are: the number of bare nuclei for C1 (barenuc, ), clump thickness for C2 (clumpthic, ), and the combination of uniformity of cell size and bland chromatin for C3 (sizeun & blandchr, ). That can be further explored in relation to the actual values and the distribution of the data entries (see below). More algorithmic details regarding EBM can also be found in Lou et al. [LCGH13].
@@
Individual and pairwise feature comparison. When a user clicks on a single feature in the bar chart (Figure 1(c), C1), a line plot and a histogram appear, displaying the impact of that specific feature (barenuc, in this case) on the model’s predictions. As illustrated in Figure 1(d), C1 (top), the line plot shows the Score, i.e., the contribution of each feature’s value inside (negative values) or outside (positive values) of the selection. It also shows the quantified uncertainty, that is, the variance in the final predictions, encoded as gray bands. For C2, for example, we can see that values of clumpthic up to 3.5 are more related to the points inside the selected cluster, with high certainty. The rest of the values are more related to the points outside the cluster, although with less certainty.
The histogram depicts the Density, which is a simple distribution of the values for a specific feature across all points globally. For C1, the feature barenuc has a clear separation between low and high values that seems to be related to the cluster formation, while for C2, clumpthic does not seem to show any clear pattern for points inside or outside the cluster (which indicates why it is important to always consider this view together with the Score line plot).
Additionally, DimVis allows users to explore the interaction effects of pairs of features, as shown in Figure 1(d). The heatmap uses color intensity to represent the level of interaction (i.e., from purple for negative Score to yellow for positive Score). In Figure 1(d), C3, we see that points where one of the two features is low while the other is high are more related to the cluster (see red boxes); when both features are low, or both are high, then the Score is higher, indicating a stronger relation to the points outside the cluster.
@@ 4 Usage Scenario: Interpreting Clusters in a Projection
Sara, a medical doctor specializing in diabetes diagnosis, recently learned about UMAP, a popular DR technique for exploring similarities in high-dimensional data. While enthusiastic about the potential of UMAP, Sara remains doubtful about the reliability of such projections. Hence, she turns to DimVis for an in-depth analysis of UMAP projections. Sara downloads the Pima Indian diabetes dataset (entries: 768; dimensions: 8; classes: 2) from the UCI ML repository [DG17] and uploads it to DimVis (DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(a.1)).
Hyperparameter adjustment. Initially, Sara tries out different hyperparameter combinations (e.g., by setting Neighbors to 10 and Min Dist to 0.4) to determine if different patterns are evident from the projections. Although these changes generate slightly different cluster shapes (not visible due to space limits), the projection still comprises a mixture of both class labels. Thus, she resets the projection to its defaults (see DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(a.2)), and the four distinct clusters combining patients Positive and Negative to diabetes appear in DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(b). Sara concludes that a more detailed, dimension-specific analysis is required to explore the underlying patterns.
Global feature importances. Sara knows that without any selection, DimVis trains a supervised EBM model (precision: ; recall: ) and visualizes the Global Feature Importances from the most to the least important: Glucose, BMI, Age, DPF, Insulin, Pregnan, SkinThick, and BloodPr (not shown due to space limits).
@@
Understanding the formation of subclusters. Now that she has gained an overview of informative dimensions, Sara wants to investigate why a cluster in DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(b) is split into left and right subclusters. By using a lasso to choose the left subcluster (cf. DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(b), lasso selection), the tool fits the EBM model (precision and recall: ; see DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(a.3)) to this selection of points (against the rest) and recomputes feature importance locally. Insulin and SkinThick are the two most important dimensions and require further examination (see DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(c)). When selecting the right subcluster, the feature ranking produces the opposite result (SkinThick first and then Insulin). She acknowledges that these two dimensions are by far the most important for the formation of this entire cluster.
@@
Commonalities between the two subclusters. Sara clicks on the first feature, which is Insulin (DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(c), red box), and the Score and Density graphs appear in DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(d.1 and d.2). In DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(d.1), negative Score, she notices that the values of the selected points for Insulin are tremendously low with most entries in the first bucket [-] (DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(d.2)). This feature follows a linear trend with high uncertainty (see the wide, gray band) for the rest of the data points, as shown in DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(d.1), positive Score. However, exactly the same effect is observable for the right subcluster (omitted due to space limits). She decides to continue with the exploration of the second important feature, which is SkinThick.
The main difference between the two subclusters. Consequently, Sara checks the individual feature importance of SkinThick for the left and right subclusters, as presented in Figure 2, (a) and (b), respectively. She understands that the former contains mostly points with moderate values (cf. Density graph) and low uncertainty for this dimension (tight gray bar in Score graph; red box), while the latter demonstrates most points concentrated in values [-] with somewhat moderate uncertainty (see the red box in Figure 2(b)). Notably, Sara identified the opposite effect for SkinThick in the right subcluster compared to what she had seen in the inspection of the left subcluster. “The common trait for this entire cluster is Insulin, while the reason behind its division into two subclusters is the different values in SkinThick,” said Sara.

### v2

# 3mm(105mm,280mm) 0

@@ Abstract
Dimensionality Reduction (DR) techniques such as t-SNE and UMAP are popular for transforming complex datasets into simpler visual representations. However, while effective in uncovering general dataset patterns, these methods may introduce artifacts and suffer from interpretability issues. This paper presents DimVis, a visualization tool that employs supervised Explainable Boosting Machine (EBM) models (trained on user-selected data of interest) as an interpretation assistant for DR projections. Our tool facilitates high-dimensional data analysis by providing an interpretation of feature relevance in visual clusters through interactive exploration of UMAP projections. Specifically, DimVis uses a contrastive EBM model that is trained in real time to differentiate between the data inside and outside a cluster of interest. Taking advantage of the inherent explainable nature of the EBM, we then use this model to interpret the cluster itself via single and pairwise feature comparisons in a ranking based on the EBM model’s feature importance. The applicability and effectiveness of DimVis are demonstrated via a use case and a usage scenario with real-world data. We also discuss the limitations and potential directions for future research.
@@ 1 Introduction
Dimensionality Reduction (DR) is an important component of the visual exploration of complex and high-dimensional datasets, transforming them into more manageable (but still representative) lower-dimensional forms with the aim to facilitate interpretation and improve computational efficiency while retaining essential information for analysis [SZS∗17]. However, the outcomes of modern nonlinear DR techniques such as t-SNE [MH08] and UMAP [MHM18] can be challenging to interpret due to the lack of inherent meaning to the extracted shapes/clusters, suboptimal hyperparameters (e.g., [WVJ16]), and misleading distortions introduced during the process [HFA17, NA18]. To address these challenges, some visualization tools have been developed with the goal of inspecting DR projections interactively to assess the global and local quality of the 2D space [CMJK20, CMJ∗20, CKK24] and to produce meaningful interpretations to the extracted shapes and clusters [FKM19, MJE21]. One recent example from Bibal et al. [BCDF21, BDF23] involved training an interpretable supervised Machine Learning (ML) model, the decision tree (DT), and use it to interpret DR layouts. However, DTs suffer from scalability issues, getting quickly deep and unparsable, and may overfit the data to achieve high fidelity. One way to overcome these limitations is by using the Explainable Boosting Machine (EBM) [LCGH13], a tree-based, cyclic gradient boosting generalized additive model, which is a powerful alternative that remains interpretable and scalable even with large-scale, high-dimensional datasets [CLG∗15].
We propose a new visualization tool, DimVis, which employs EBM models for unveiling the factors influencing cluster formation and data relationships in nonlinear projections. It enables users to interactively inspect visual patterns of the DR layout (clusters, shapes, etc.) to gain insights about their content by training an EBM model on-the-fly on user-selected data points. The explanation is based on exploring single and pairs of features ranked by EBM’s intrinsic feature importance regarding their ability to explain the separation between the selection and the remaining points of the dataset. The results are visualized with intuitive and simple visualizations such as bar charts, line plots, histograms, and heatmaps. We demonstrate DimVis via a use case and a usage scenario with healthcare datasets (also used, for example, in [MQB19] and [CMK20]), showcasing the tool’s capability of analyzing high-dimensional datasets and enhancing the interpretability of the resulting 2D space from a DR technique.
@@ 2 Related Work
In early research on DR interpretation, static visualizations were generated by considering the layout as a whole, without giving the users the ability and flexibility to specify their own regions of interest [dSRM∗15]. Interactive interpretation of selections and groups, as is the case for DimVis, has been explored for example in Probing Projections [SDMT16], where groups of points (obtained either by selection or a clustering algorithm) are compared to each other or to the rest of the dataset via individual histograms for each dimension; the users then must visually determine dimension importance by themselves. This approach provides a strong foundation but may present problems with very high-dimensional datasets, and the interpretations are inherently univariate.
Some examples of recent work that are more related to DimVis are those from Marcilio-Jr et al. [MJE21] and Fujiwara et al. [FKM19]. In both cases, users can define groups of points of interest and interpret them by contrasting their features against outside points. Fujiwara et al. use an adapted version of contrastive PCA [AZBZ18], which in itself is an adapted version of the classic PCA for highlighting the differences between two sets of data points. Although relevant, we argue that being a linear algorithm, interpretations based on adapted versions of PCA will still have limitations when dealing with very high-dimensional and complex groups of points. Marcilio-Jr et al. [MJE21] used Shapley values to generate explanations for each data point based on its distance to the centroids of the specified clusters, which are then visualized with a relatively complex visual abstraction based on merging histograms and connecting them with edges (as lines). In contrast, DimVis uses basic, simple visualizations that are arguably more natural and easily interpreted, and the explanations come directly from an inherently interpretable and robust supervised model instead of being generated with a black-box, model-agnostic method.
DT-SNE [BDF23] is a simpler but static approach combining t-SNE with supervised decision trees for interpreting projections with features, and IXVC [BCDF21] supports users in interactively explaining projections with decision trees. However, decision trees, especially visualized as node-link diagrams, can become quickly hard to interpret, and the increase in the number of dataset features may heavily affect their scalability. On the contrary, DimVis uses EBM – which can be as powerful as ensemble learning models – and offers simple and scalable visual representations for exploring the most important features in user-selected data contrastively.
@@ 3 DimVis: System Overview and Use Case
DimVis is an open-source, web-based visualization tool [Dim23] which uses the state-of-the-art, supervised “glass-box” EBM model [LCGH13] to interpret visualizations generated with unsupervised DR techniques. We chose the EBM because its results and pre-defined visualizations (which we used) have already been successfully evaluated with real-world domain experts [CLG∗15].
After selecting the dataset (shown in DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(a.1)), initially only views indicated in (a.1), (a.2), and (b) are shown. To initiate the exploration and trigger the remaining views, a lasso selection of data points must be made in the UMAP view (b) containing a visual cluster perceived by the user as being of interest. As soon as a selection is made, an EBM model is trained in the background (more details below) and the remaining visualizations are presented. The model’s precision and recall are visible (in DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(a.3)), as well as single and pairwise feature importances extracted from the EBM model and visualized in a bar chart sorted in descending order from top to bottom (see DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(c)). On click, users can explore further a feature with a Score line plot (DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(d.1)) and a Density histogram (DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(d.2)), or two features with a Score heatmap.
@@
DimVis utilizes the UMAP algorithm [MHM18] and users can interactively adjust UMAP’s hyperparameters – “Number of Neighbors” (Neighbors) and “Minimum Distance” (Min Dist) – to explore different projections (see Figure 1(a)). The projection in Figure 1(b) contains three distinct clusters (C1–C3), and no GT labels to simulate a typical unsupervised learning scenario. When a group of points is selected with the lasso functionality, the EBM gets retrained and selected points are marked with class while the remaining points are marked with class . Using the inherent interpretability features of the EBM model, we rank the main features (single and pairs) that contribute to the separation of the cluster according to the trained model, using the Mean Absolute Score (Weighted). As shown in Figure 1(c), the most important single and pair of features are: the number of bare nuclei for C1 (barenuc, ), clump thickness for C2 (clumpthic, ), and the combination of uniformity of cell size and bland chromatin for C3 (sizeun & blandchr, ). That can be further explored in relation to the actual values and the distribution of the data entries (see below). More algorithmic details regarding EBM can also be found in Lou et al. [LCGH13].
@@
When a user clicks on a single feature in the bar chart (Figure 1(c), C1), a line plot and a histogram appear, displaying the impact of that specific feature (barenuc, in this case) on the model’s predictions. As illustrated in Figure 1(d), C1 (top), the line plot shows the Score, i.e., the contribution of each feature’s value inside (negative values) or outside (positive values) of the selection. It also shows the quantified uncertainty, that is, the variance in the final predictions, encoded as gray bands. For C2, for example, we can see that values of clumpthic up to 3.5 are more related to the points inside the selected cluster, with high certainty. The rest of the values are more related to the points outside the cluster, although with less certainty. The histogram depicts the Density, which is a simple distribution of the values for a specific feature across all points globally. For C1, the feature barenuc has a clear separation between low and high values that seems to be related to the cluster formation, while for C2, clumpthic does not seem to show any clear pattern for points inside or outside the cluster (which shows why it is important to consider this view together with the Score line plot). Additionally, DimVis allows users to explore pairs of features, as shown in Figure 1(d). The heatmap uses color intensity to represent the level of interaction (i.e., from purple for negative Score to yellow for positive Score). In Figure 1(d), C3, we see that points where one of the two features is low while the other is high are more related to the cluster (see green boxes); when both features are low, or both are high, then the Score is higher, indicating a stronger relation to the points outside the cluster.
@@ 4 Usage Scenario: Interpreting Clusters in a Projection
For this hypothetical scenario, we follow a physician specializing in diabetes (which we will call Sara). She recently learned about UMAP but, while enthusiastic, she remains doubtful about the reliability of such projections. Hence, she turns to DimVis for an in-depth analysis of UMAP projections. Sara uploads to DimVis (DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(a.1)) the Pima Indian diabetes dataset (entries: 768; dimensions: 8; classes: 2) from the UCI ML repository [DG17].
@@
Without any selection, DimVis trains a supervised EBM model (precision: ; recall: ) and visualizes the Global Feature Importances from the most to the least important: Glucose, BMI, Age, DPF, Insulin, Pregnan, SkinThick, and BloodPr (not shown due to space limits). Now that she has gained an overview, Sara wants to investigate why a cluster in DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(b) is split into left and right subclusters. By using a lasso to choose the left subcluster (cf. DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(b)), the tool fits the EBM model (precision and recall: ; see DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(a.3)) to this selection of points (against the rest) and recomputes feature importance locally. Insulin and SkinThick are the two most important dimensions and require further examination (see DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(c)). When selecting the right subcluster, the feature ranking produces the opposite result (SkinThick first and then Insulin). She acknowledges that these two dimensions are by far the most important for the formation of this entire cluster.
@@
Sara clicks on the first feature, which is Insulin (DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(c), green box), and the Score and Density graphs appear in DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(d.1 and d.2). In DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(d.1), negative Score, she notices that the values of the selected points for Insulin are very low with most entries in the first bucket [-] (DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(d.2)). This feature follows a linear trend with high uncertainty (see the wide, gray band) for the rest of the data points, as shown in DimVis: Interpreting Visual Clusters in Dimensionality Reduction With Explainable Boosting Machine(d.1), positive Score. However, the same effect is observable for the right subcluster (omitted due to space limits). She decides to continue with the exploration of the second important feature, which is SkinThick. She checks the individual feature importance of SkinThick for the left and right subclusters, as presented in Figure 2, (a) and (b), respectively. She understands that the former contains mostly points with moderate values (cf. Density graph) and low uncertainty for this dimension (tight gray bar in Score graph; green box), while the latter demonstrates most points concentrated in values [-] with somewhat moderate uncertainty (see the green box in Figure 2(b)). Notably, Sara identified the opposite effect for SkinThick in the right subcluster compared to what she had seen in the inspection of the left subcluster. She concludes, then, that the common trait for this entire cluster is Insulin, while the reason behind its division into two subclusters is the different values in SkinThick.